Buzz's Educational Software Outlet sells two or more of the video games as a single package. Managers are keenly interested in individual product−profitability figures. Information pertaining to three bundled products and the stand−alone prices is as​ follows:

Stand−Alone Selling Price

Cost

Package

Packaged Price

Reading Fun

$16

$3.40

1. Reading Fun​ & Math Fun

$42

Math Fun

$36

$4.05

2. Reading Fun​ & Analysis

$50

Analysis

$48

$5.45

3. All three

$76

Using the stand−alone method with selling price as the weight for revenue​ allocation, what amount of revenue will be allocated to Math Fun in the package that contains all three​ products? (Do not round any intermediary​ calculations.)

A. $32.57

B. $18.00

C. $36.00

D. $27.36
Answer #1
Stand−Alone Selling Price Percentage of Stand−Alone Selling Price Selling price allocated
Reading Fun $16 16/100 = 16% 76 x 16% = 12.16
Math Fun $36 36/100 = 36% 76 x 36% = 27.36
Analysis $48 48/100 = 48% 76 x 48% = 36.48
Total $100 100% $76

Revenue allocated to Math fun = $27.36

Correct option is (D)

kindly give a positive rating if you are satisfied with the solution. Please ask if you have any query related to the question, Thanks.